OpenStudy (anonymous):

how to find the exact value of sin 200 degrees

OpenStudy (anonymous):

sin(200)=-sin(20) sin(200)=sin(340)

OpenStudy (anonymous):

I see how it is same as sin 20 but how do i find the exact value

OpenStudy (anonymous):

I don't know how to do but you'll use Taylor Series.

Parth (parthkohli):

\[\rm f(x) = \sum_{n = 0}^{\infty} {f^{(n)}(0)\over n!}x^n\]My favorite Maclaurin Series.